dy Find the equation of the curve that passes through (-1, -3) if its slope is given by 18x2 - 8x for each x ya
Answer #1

dy dx 18.x-8x dr = (18-28)Jk 18. 3 - ४ 8.2 = B 2 + 3 2 ४ = 6x3-५x+c given when X=-13 43-3 -3 = 3 6 (-1)3-५ (-1)+C -3 = -6-५ +